# Behind Closed Words: Creating and Investigating the forePLay Annotated Dataset for Polish Erotic Discourse Warning: This paper contains examples of sexually explicit material

### v1

@@ 5.3 Polish Large Language Models
Our evaluation framework encompasses three variants from the PLLuM (Polish Large Language Model) family: PLLuM-Llama 3.1 8B-Instruct, PLLuM-Mistral 12B, and PLLuM-Mixtral-8x7B, adapted from their respective base architectures. These models underwent pre-training on 147B Polish language tokens and subsequent instruction tuning. We additionally employed Bielik 2.3, trained on 200B tokens and aligned using DPO-positive on 66,000 examples. To assess adaptation capabilities, we fine-tuned PLLuM-Llama 3.1 8B and PLLuM-Mistral 12B on our classification task. All base models were evaluated under 0-shot, 1-shot, and 5-shot configurations, while fine-tuned variants underwent standard supervised evaluation.
@@ 6 Results
Bielik-11B-v2.2-Instruct 0.848 0.858 0.868 0.601 0.597 0.607 0.427 0.416 0.423 0.312 0.437 0.48
PLLuM-Llama-3.1-8B 0.641 0.737 0.709 0.425 0.446 0.425 0.301 0.324 0.326 0.314 0.274 0.273
PLLuM-Mistral 12B 0.894 0.842 0.844 0.656 0.610 0.613 0.456 0.441 0.427 0.401 0.359 0.401
PLLuM-Mixtral 8x7B 0.874 0.902 0.921 0.647 0.662 0.670 0.452 0.456 0.426 0.422 0.406 0.422
PLLuM-Llama-3.1-8B (SFT) 0.947 0.764 0.461 0.435
PLLuM-Mistral 12B (SFT) 0.946 0.792 0.458 0.488
In the case of LLMs, models from the PLLuM family consistently demonstrate strong performance across varying label counts and shot configurations, with particular emphasis on the PLLuM-Mixtral 8x7B, which achieves outstanding F1 scores across a wide range of label counts, notably reaching 0.921 for Basic data and 0.670 for Core data in the 5-shot setting. Additionally PLLuM-Llama-3.1-8B (SFT) and PLLuM-Mistral 12B (SFT) show significant improvements when subjected to supervised fine-tuning, resulting in enhanced performance, with the highest F1 scores observed across various label counts.
@@
PLLuM-Mistral 12B 3 5 1 3 2 3 8 7 0 6 10 5 53
PLLuM-Mixtral 8x7B 0 0 1 0 0 2 0 0 2 0 1 2 8
PLLuM-Llama-3.1-8B 1 4 0 1 4 3 1 3 5 0 4 1 27
PLLuM-Mistral 12B (SFT) 0 0 0 0 0
PLLuM-Llama-3.1-8B (SFT) 0 0 0 0 0
@@ 7 Discussion
The PLLuM family of models exhibits exceptional few-shot learning capabilities, with PLLuM-Mixtral 8x7B achieving F1 scores of $0.921$ in 5-shot settings for binary classification, while supervised fine-tuning of PLLuM-Mistral 12B yields the highest overall performance (F1=$0.946$). However, the diminishing returns observed with increasing model size, coupled with the persistent challenge of handling ambiguous content in imbalanced datasets, suggests that architectural sophistication alone may not overcome the fundamental challenges of erotic content detection in morphologically complex languages. These findings underscore the importance of language-specific model development for sensitive content detection tasks, rather than relying solely on increasingly large multilingual models.
@@ Appendix D Computational Details
The encoder-based models were trained for approximately 5 hours and 10 hours, respectively, for the Base and Large versions, using A100 (40GB) GPU with early stopping. The batch size was set to 8 for both model variants. Training of the PLLuM family models on our dataset took approximately 3 hours for the PLLuM-Llama 3.1 8B (SFT) model and 4 hours for the PLLuM-Mistral 12B (SFT) model. These models were trained on a distributed setup consisting of 2 nodes, each equipped with 4 H100 (96GB) GPUs, ensuring parallel processing across GPUs for enhanced training efficiency.

### v2

@@ 5.3 Polish Large Language Models
Our evaluation framework encompasses three variants from the PLLuM (Polish Large Language Model) family: Llama-3.1-8B-PLLuM, PLLuM-Mistral-12B, and PLLuM-Mixtral-8x7B, adapted from their respective base architectures. These models underwent pre-training on 147B Polish language tokens and subsequent instruction tuning. We additionally employed Bielik 2.3, trained on 200B tokens and aligned using DPO-positive on 66,000 examples. To assess adaptation capabilities, we fine-tuned Llama-3.1-8B-PLLuM and PLLuM-Mistral-12B on our classification task. All base models were evaluated under 0-shot, 1-shot, and 5-shot configurations, while fine-tuned variants underwent standard supervised evaluation.
@@ 6 Results
Bielik-11B-v2.3-Instruct 0.848 0.858 0.868 0.601 0.597 0.607 0.427 0.416 0.423 0.312 0.437 0.48
Llama-3.1-8B-PLLuM 0.641 0.737 0.709 0.425 0.446 0.425 0.301 0.324 0.326 0.314 0.274 0.273
PLLuM-Mistral-12B 0.894 0.842 0.844 0.656 0.610 0.613 0.456 0.441 0.427 0.401 0.359 0.401
PLLuM-Mixtral-8x7B 0.874 0.902 0.921 0.647 0.662 0.670 0.452 0.456 0.426 0.422 0.406 0.422
Llama-3.1-8B-PLLuM (SFT) 0.947 0.764 0.461 0.435
PLLuM-Mistral-12B (SFT) 0.946 0.792 0.458 0.488
In the case of LLMs, models from the PLLuM family consistently demonstrate strong performance across varying label counts and shot configurations, with particular emphasis on the PLLuM-Mixtral-8x7B, which achieves outstanding F1 scores across a wide range of label counts, notably reaching 0.921 for Basic data and 0.670 for Core data in the 5-shot setting. Additionally Llama-3.1-8B-PLLuM (SFT) and PLLuM-Mistral-12B (SFT) show significant improvements when subjected to supervised fine-tuning, resulting in enhanced performance, with the highest F1 scores observed across various label counts.
@@
PLLuM-Mistral-12B 3 5 1 3 2 3 8 7 0 6 10 5 53
PLLuM-Mixtral-8x7B 0 0 1 0 0 2 0 0 2 0 1 2 8
Llama-3.1-8B-PLLuM 1 4 0 1 4 3 1 3 5 0 4 1 27
PLLuM-Mistral-12B (SFT) 0 0 0 0 0
Llama-3.1-8B-PLLuM (SFT) 0 0 0 0 0
@@ 7 Discussion
The PLLuM family of models exhibits exceptional few-shot learning capabilities, with PLLuM-Mixtral-8x7B achieving F1 scores of $0.921$ in 5-shot settings for binary classification, while supervised fine-tuning of PLLuM-Mistral-12B yields the highest overall performance (F1=$0.946$). However, the diminishing returns observed with increasing model size, coupled with the persistent challenge of handling ambiguous content in imbalanced datasets, suggests that architectural sophistication alone may not overcome the fundamental challenges of erotic content detection in morphologically complex languages. These findings underscore the importance of language-specific model development for sensitive content detection tasks, rather than relying solely on increasingly large multilingual models.
@@ Appendix D Computational Details
The encoder-based models were trained for approximately 5 hours and 10 hours, respectively, for the Base and Large versions, using A100 (40GB) GPU with early stopping. The batch size was set to 8 for both model variants. Training of the PLLuM family models on our dataset took approximately 3 hours for the Llama-3.1-8B-PLLuM (SFT) model and 4 hours for the PLLuM-Mistral-12B (SFT) model. These models were trained on a distributed setup consisting of 2 nodes, each equipped with 4 H100 (96GB) GPUs, ensuring parallel processing across GPUs for enhanced training efficiency.